# The theory of implicit operations

### v1

@@ 10. Primitive positive expansions
Therefore, it only remains to prove that if $f\in\mathsf{ext}_{pp}(\mathsf{M})$, then $f_{*}\in\mathsf{ext}_{pp}(\mathsf{K})$. Accordingly, suppose that $f\in\mathsf{ext}_{pp}(\mathsf{M})$. Since we already proved that $f_{*}\in\mathsf{imp}_{pp}(\mathsf{K})$, it suffices to show that $f_{*}\in\mathsf{ext}(\mathsf{K})$. To this end, consider $\bm{A}\in\mathsf{K}$ and $a_{1},\dots,a_{n}\in A$. As $\mathsf{K}$ is the class of ${L}_{\mathsf{K}}$-subreducts of $\mathsf{M}$ by Proposition 10.2, there exists $\bm{B}\in\mathsf{M}$ such that $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Since $f\in\mathsf{ext}(\mathsf{M})$, $\bm{B}\in\mathsf{M}$, and $a_{1},\dots,a_{n}\in A\subseteq B$, there exists $\bm{C}\in\mathsf{M}$ such that $\bm{B}\leqslant\bm{C}$ and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{C}})$. From $\bm{C}\in\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ it follows that there also exists $\bm{D}\in\mathsf{K}[{L}_{\mathcal{F}}]$ with $\bm{C}\leqslant\bm{D}$. Since $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$ and $\bm{B}\leqslant\bm{C}\leqslant\bm{D}\in\mathsf{M}$ and $\mathsf{K}$ is the class of ${L}_{\mathsf{K}}$-subreducts of $\mathsf{M}$, we have $\bm{A}\leqslant\bm{D}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$. Moreover, by applying Proposition 8.1 to $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{C}})$ and $\bm{C}\leqslant\bm{D}$, we obtain $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{D}})$. Hence,
@@
Let $\mathsf{M}_{2}$ be a pp expansion of a pp expansion $\mathsf{M}_{1}$ of $\mathsf{K}$. We will prove that $\mathsf{M}_{2}$ is also a pp expansion of $\mathsf{K}$. First, as $\mathsf{M}_{1}$ is a pp expansion of $\mathsf{K}$, there exist $\mathcal{F}_{1}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ and an $\mathcal{F}_{1}$-expansion ${L}_{\mathcal{F}_{1}}$ of ${L}_{\mathsf{K}}$ such that $\mathsf{M}_{1}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}_{1}}])$. Similarly, as $\mathsf{M}_{2}$ is a pp expansion of $\mathsf{M}_{1}$, there exist $\mathcal{F}_{2}\subseteq\mathsf{ext}_{pp}(\mathsf{M}_{1})$ and an $\mathcal{F}_{2}$-expansion ${L}_{\mathcal{F}_{2}}$ of ${L}_{\mathcal{F}_{1}}$ such that $\mathsf{M}_{2}=\mathbb{S}(\mathsf{M}_{1}[{L}_{\mathcal{F}_{2}}])$. Hence,
@@
As $\mathcal{F}_{2}\subseteq\mathsf{ext}_{pp}(\mathsf{M}_{1})$ and $\mathsf{M}_{1}$ is a pp expansion of $\mathsf{K}$, Proposition 10.22(iii) guarantees that for every $f\in\mathcal{F}_{2}$ there exists $f_{*}\in\mathsf{ext}_{pp}(\mathsf{K})$ such that
@@
Since $\mathcal{F}_{1}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ by assumption, the set
@@
is a subset of $\mathsf{ext}_{pp}(\mathsf{K})$. Define $g_{f_{*}}=g_{f}$ for each $f\in\mathcal{F}_{2}$ and consider the following $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$:
@@
The equalities above are justified as follows. To prove the first, recall that $\bm{A}\in\mathsf{M}_{1}[{L}_{\mathcal{F}_{2}}]$ and $f\in\mathcal{F}_{2}$, whence $f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}}$ is a total function and $g_{f}^{\bm{A}}(a_{1},\dots,a_{n})=f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}}(a_{1},\dots,a_{n})$. To prove the second, observe that the assumption that $f\in\mathcal{F}_{2}\subseteq\mathsf{imp}(\mathsf{M}_{1})$, $\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}\leqslant\bm{B}$, and $\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}},\bm{B}\in\mathsf{M}_{1}$ allows us to apply 8.1 to the fact that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}})$, obtaining $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})$ and $f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}}(a_{1},\dots,a_{n})=f^{\bm{B}}(a_{1},\dots,a_{n})$. The third equality follows from $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}_{1}}]$ and (46). To prove the fourth, observe that $f_{*}\in\mathsf{imp}(\mathsf{K})$, $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$, and $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Therefore, we can apply 8.1 to the fact that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f_{*}^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}})$, obtaining $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f_{*}^{\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}})$ and $f_{*}^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a_{1},\dots,a_{n})=f_{*}^{\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a_{1},\dots,a_{n})$. The fifth holds because $\bm{C}\in\mathsf{K}[{L}_{\mathcal{F}}]$ and $f_{*}\in\mathcal{F}$. Lastly, the sixth equality holds because $g_{f_{*}}=g_{f}$ by assumption. This concludes the proof of (48). Hence, we obtain $\bm{A}\leqslant\bm{C}$, as desired. ∎
@@
The proof of 10.7 hinges on the following result, which describes how to lift implicit operations to pp expansions.
@@
    if $f\in\mathsf{imp}_{pp}(\mathsf{K})$, then $f_{\bullet}\in\mathsf{imp}_{pp}(\mathsf{M})$;
@@
(i): Observe that 10.2 implies that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$ for every $\bm{A}\in\mathsf{M}$. Hence, $f_{\bullet}$ is a well-defined partial function on $\mathsf{M}$. We show that $f_{\bullet}$ is preserved by homomorphisms in $\mathsf{M}$. Consider a homomorphism $h\colon\bm{A}\to\bm{B}$ with $\bm{A},\bm{B}\in\mathsf{M}$ and let $a_{1},\dots,a_{n}\in A$ be such that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f_{\bullet}^{\bm{A}})$. Since $f_{\bullet}^{\bm{A}}=f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$, we have
@@
(ii): Suppose that $f\in\mathsf{imp}_{pp}(\mathsf{K})$. Then we may assume that $\varphi$ is a pp formula. It follows from (i) that $f_{\bullet}$ is defined by $\varphi$ and belongs to $\mathsf{imp}(\mathsf{M})$, whence $f_{\bullet}\in\mathsf{imp}_{pp}(\mathsf{M})$.
(iii): Since $\mathsf{M}$ is a pp expansion of $\mathsf{K}$, it is of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ for some $\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ and ${L}_{\mathcal{F}}$. Suppose that $f\in\mathsf{ext}(\mathsf{K})$. To show that $f_{\bullet}\in\mathsf{ext}(\mathsf{M})$, consider $\bm{A}\in\mathsf{M}$ and $a_{1},\dots,a_{n}\in A$. Since $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, there exists $\bm{B}[{L}_{\mathcal{F}}]\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{A}\leqslant\bm{B}[{L}_{\mathcal{F}}]$. As $\bm{B}\in\mathsf{K}$, 8.4 yields $\bm{C}\in\mathsf{K}$ such that $\bm{B}\leqslant\bm{C}$ and $g^{\bm{C}}$ is total for every $g\in\mathsf{ext}(\mathsf{K})$. In particular,
@@
By 9.5 from $\bm{B}\leqslant\bm{C}$ it follows that $\bm{B}[{L}_{\mathcal{F}}]\leqslant\bm{C}[{L}_{\mathcal{F}}]$. Then $\bm{A}\leqslant\bm{B}[{L}_{\mathcal{F}}]\leqslant\bm{C}[{L}_{\mathcal{F}}]$ with $\bm{C}[{L}_{\mathcal{F}}]\in\mathsf{K}[{L}_{\mathcal{F}}]\subseteq\mathsf{M}$. Since $\bm{C}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}=\bm{C}$, the definition of $f_{\bullet}$ yields $f_{\bullet}^{\bm{C}[{L}_{\mathcal{F}}]}=f^{\bm{C}}$. From (51) it follows that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f_{\bullet}^{\bm{C}[{L}_{\mathcal{F}}]})$ and, therefore, $f_{\bullet}\in\mathsf{ext}(\mathsf{M})$. ∎
We now proceed to prove 10.7.
@@
Let $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ and ${L}_{\mathcal{G}}={L}_{\mathsf{K}}\cup\{g_{f}:f\in\mathcal{G}\}$. Define $\mathcal{G}_{\bullet}=\{f_{\bullet}:f\in\mathcal{G}-\mathcal{F}\}$, where $f_{\bullet}=\langle f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}:\bm{A}\in\mathsf{M}\rangle$. Since $\mathcal{G}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$, 10.26 implies that $\mathcal{G}_{\bullet}\subseteq\mathsf{ext}_{pp}(\mathsf{M})$. Consider the pp expansion $\mathbb{S}(\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}])$ of $\mathsf{M}$ induced by $\mathcal{G}_{\bullet}$ and ${L}_{\mathcal{G}_{\bullet}}$, where ${L}_{\mathcal{G}_{\bullet}}={L}_{\mathcal{G}}$ and $g_{f}^{\bm{A}[{L}_{\mathcal{G}_{\bullet}}]}=f_{\bullet}^{\bm{A}}$ for all $f\in\mathcal{G}$-$\mathcal{F}$ and $\bm{A}[{L}_{\mathcal{G}_{\bullet}}]\in\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}]$. Since $\mathbb{S}(\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}])$ is a pp expansion of $\mathsf{M}$, to prove that $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{G}}])$ is a pp expansion of $\mathsf{M}$, it suffices to show that
@@
For the inclusion from left to right of (52) it is enough to show that $\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}]\subseteq\mathbb{S}(\mathsf{K}[{L}_{\mathcal{G}}])$. To this end, consider $\bm{A}\in\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}]$. Then $\bm{A}{\upharpoonright}_{{L}_{\mathcal{F}}}\in\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. Consequently, there exists $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{A}{\upharpoonright}_{{L}_{\mathcal{F}}}\leqslant\bm{B}$. As $\mathcal{G}\subseteq\mathsf{ext}(\mathsf{K})$, 9.6 yields that $\mathsf{K}$ is the class of subreducts of $\mathsf{K}[{L}_{\mathcal{G}}]$. Together with $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$ (which holds because $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$), this entails that there exists $\bm{C}[{L}_{\mathcal{G}}]\in\mathsf{K}[{L}_{\mathcal{G}}]$ such that $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{C}$.
@@
The first equality above holds because $\bm{A}\in\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}]$ and the second is a consequence of the definition of $f_{\bullet}$. The third equality follows from 8.1 because $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{C}$, and the last from the interpretation of $g_{f}$ in $\mathsf{K}[{L}_{\mathcal{G}}]$. Therefore, $\bm{A}\leqslant\bm{C}[{L}_{\mathcal{G}}]\in\mathsf{K}[{L}_{\mathcal{G}}]$. We conclude that $\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}]\subseteq\mathbb{S}(\mathsf{K}[{L}_{\mathcal{G}}])$, as desired.

### v2

@@ 10. Primitive positive expansions
Therefore, it only remains to prove that if $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{M})$, then $f_{*}\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$. Accordingly, suppose that $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{M})$. Since we already proved that $f_{*}\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$, it suffices to show that $f_{*}\in\mathsf{ext}(\mathsf{K})$. To this end, consider $\bm{A}\in\mathsf{K}$ and $a_{1},\dots,a_{n}\in A$. As $\mathsf{K}$ is the class of ${L}_{\mathsf{K}}$-subreducts of $\mathsf{M}$ by Proposition 10.2, there exists $\bm{B}\in\mathsf{M}$ such that $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Since $f\in\mathsf{ext}(\mathsf{M})$, $\bm{B}\in\mathsf{M}$, and $a_{1},\dots,a_{n}\in A\subseteq B$, there exists $\bm{C}\in\mathsf{M}$ such that $\bm{B}\leqslant\bm{C}$ and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{C}})$. From $\bm{C}\in\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ it follows that there also exists $\bm{D}\in\mathsf{K}[{L}_{\mathcal{F}}]$ with $\bm{C}\leqslant\bm{D}$. Since $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$ and $\bm{B}\leqslant\bm{C}\leqslant\bm{D}\in\mathsf{M}$ and $\mathsf{K}$ is the class of ${L}_{\mathsf{K}}$-subreducts of $\mathsf{M}$, we have $\bm{A}\leqslant\bm{D}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$. Moreover, by applying Proposition 8.1 to $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{C}})$ and $\bm{C}\leqslant\bm{D}$, we obtain $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{D}})$. Hence,
@@
Let $\mathsf{M}_{2}$ be a pp expansion of a pp expansion $\mathsf{M}_{1}$ of $\mathsf{K}$. We will prove that $\mathsf{M}_{2}$ is also a pp expansion of $\mathsf{K}$. First, as $\mathsf{M}_{1}$ is a pp expansion of $\mathsf{K}$, there exist $\mathcal{F}_{1}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ and an $\mathcal{F}_{1}$-expansion ${L}_{\mathcal{F}_{1}}$ of ${L}_{\mathsf{K}}$ such that $\mathsf{M}_{1}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}_{1}}])$. Similarly, as $\mathsf{M}_{2}$ is a pp expansion of $\mathsf{M}_{1}$, there exist $\mathcal{F}_{2}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{M}_{1})$ and an $\mathcal{F}_{2}$-expansion ${L}_{\mathcal{F}_{2}}$ of ${L}_{\mathcal{F}_{1}}$ such that $\mathsf{M}_{2}=\mathbb{S}(\mathsf{M}_{1}[{L}_{\mathcal{F}_{2}}])$. Hence,
@@
As $\mathcal{F}_{2}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{M}_{1})$ and $\mathsf{M}_{1}$ is a pp expansion of $\mathsf{K}$, Proposition 10.22(iii) guarantees that for every $f\in\mathcal{F}_{2}$ there exists $f_{*}\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ such that
@@
Since $\mathcal{F}_{1}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ by assumption, the set
@@
is a subset of $\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$. Define $g_{f_{*}}=g_{f}$ for each $f\in\mathcal{F}_{2}$ and consider the following $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$:
@@
The equalities above are justified as follows. To prove the first, recall that $\bm{A}\in\mathsf{M}_{1}[{L}_{\mathcal{F}_{2}}]$ and $f\in\mathcal{F}_{2}$, whence $f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}}$ is a total function and $g_{f}^{\bm{A}}(a_{1},\dots,a_{n})=f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}}(a_{1},\dots,a_{n})$. To prove the second, observe that the assumption that $f\in\mathcal{F}_{2}\subseteq\mathsf{imp}(\mathsf{M}_{1})$, $\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}\leqslant\bm{B}$, and $\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}},\bm{B}\in\mathsf{M}_{1}$ allows us to apply Proposition 8.1 to the fact that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}})$, obtaining $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})$ and $f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}}(a_{1},\dots,a_{n})=f^{\bm{B}}(a_{1},\dots,a_{n})$. The third equality follows from $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}_{1}}]$ and (46). To prove the fourth, observe that $f_{*}\in\mathsf{imp}(\mathsf{K})$, $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$, and $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Therefore, we can apply Proposition 8.1 to the fact that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f_{*}^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}})$, obtaining $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f_{*}^{\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}})$ and $f_{*}^{\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a_{1},\dots,a_{n})=f_{*}^{\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a_{1},\dots,a_{n})$. The fifth holds because $\bm{C}\in\mathsf{K}[{L}_{\mathcal{F}}]$ and $f_{*}\in\mathcal{F}$. Lastly, the sixth equality holds because $g_{f_{*}}=g_{f}$ by assumption. This concludes the proof of (48). Hence, we obtain $\bm{A}\leqslant\bm{C}$, as desired. ∎
@@
The proof of Theorem 10.7 hinges on the following result, which describes how to lift implicit operations to pp expansions.
@@
    if $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$, then $f_{\bullet}\in\mathsf{imp}_{\textsc{pp}}(\mathsf{M})$;
@@
(i): Observe that Proposition 10.2 implies that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$ for every $\bm{A}\in\mathsf{M}$. Hence, $f_{\bullet}$ is a well-defined partial function on $\mathsf{M}$. We show that $f_{\bullet}$ is preserved by homomorphisms in $\mathsf{M}$. Consider a homomorphism $h\colon\bm{A}\to\bm{B}$ with $\bm{A},\bm{B}\in\mathsf{M}$ and let $a_{1},\dots,a_{n}\in A$ be such that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f_{\bullet}^{\bm{A}})$. Since $f_{\bullet}^{\bm{A}}=f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$, we have
@@
(ii): Suppose that $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$. Then we may assume that $\varphi$ is a pp formula. It follows from (i) that $f_{\bullet}$ is defined by $\varphi$ and belongs to $\mathsf{imp}(\mathsf{M})$, whence $f_{\bullet}\in\mathsf{imp}_{\textsc{pp}}(\mathsf{M})$.
(iii): Since $\mathsf{M}$ is a pp expansion of $\mathsf{K}$, it is of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ for some $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ and ${L}_{\mathcal{F}}$. Suppose that $f\in\mathsf{ext}(\mathsf{K})$. To show that $f_{\bullet}\in\mathsf{ext}(\mathsf{M})$, consider $\bm{A}\in\mathsf{M}$ and $a_{1},\dots,a_{n}\in A$. Since $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, there exists $\bm{B}[{L}_{\mathcal{F}}]\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{A}\leqslant\bm{B}[{L}_{\mathcal{F}}]$. As $\bm{B}\in\mathsf{K}$, Theorem 8.4 yields $\bm{C}\in\mathsf{K}$ such that $\bm{B}\leqslant\bm{C}$ and $g^{\bm{C}}$ is total for every $g\in\mathsf{ext}(\mathsf{K})$. In particular,
@@
By Proposition 9.5 from $\bm{B}\leqslant\bm{C}$ it follows that $\bm{B}[{L}_{\mathcal{F}}]\leqslant\bm{C}[{L}_{\mathcal{F}}]$. Then $\bm{A}\leqslant\bm{B}[{L}_{\mathcal{F}}]\leqslant\bm{C}[{L}_{\mathcal{F}}]$ with $\bm{C}[{L}_{\mathcal{F}}]\in\mathsf{K}[{L}_{\mathcal{F}}]\subseteq\mathsf{M}$. Since $\bm{C}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}=\bm{C}$, the definition of $f_{\bullet}$ yields $f_{\bullet}^{\bm{C}[{L}_{\mathcal{F}}]}=f^{\bm{C}}$. From (51) it follows that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f_{\bullet}^{\bm{C}[{L}_{\mathcal{F}}]})$ and, therefore, $f_{\bullet}\in\mathsf{ext}(\mathsf{M})$. ∎
We now proceed to prove Theorem 10.7.
@@
Let $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ and ${L}_{\mathcal{G}}={L}_{\mathsf{K}}\cup\{g_{f}:f\in\mathcal{G}\}$. Define $\mathcal{G}_{\bullet}=\{f_{\bullet}:f\in\mathcal{G}-\mathcal{F}\}$, where $f_{\bullet}=\langle f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}:\bm{A}\in\mathsf{M}\rangle$. Since $\mathcal{G}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$, Proposition 10.26 implies that $\mathcal{G}_{\bullet}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{M})$. Consider the pp expansion $\mathbb{S}(\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}])$ of $\mathsf{M}$ induced by $\mathcal{G}_{\bullet}$ and ${L}_{\mathcal{G}_{\bullet}}$, where ${L}_{\mathcal{G}_{\bullet}}={L}_{\mathcal{G}}$ and $g_{f}^{\bm{A}[{L}_{\mathcal{G}_{\bullet}}]}=f_{\bullet}^{\bm{A}}$ for all $f\in\mathcal{G}$-$\mathcal{F}$ and $\bm{A}[{L}_{\mathcal{G}_{\bullet}}]\in\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}]$. Since $\mathbb{S}(\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}])$ is a pp expansion of $\mathsf{M}$, to prove that $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{G}}])$ is a pp expansion of $\mathsf{M}$, it suffices to show that
@@
For the inclusion from left to right of (52) it is enough to show that $\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}]\subseteq\mathbb{S}(\mathsf{K}[{L}_{\mathcal{G}}])$. To this end, consider $\bm{A}\in\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}]$. Then $\bm{A}{\upharpoonright}_{{L}_{\mathcal{F}}}\in\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. Consequently, there exists $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{A}{\upharpoonright}_{{L}_{\mathcal{F}}}\leqslant\bm{B}$. As $\mathcal{G}\subseteq\mathsf{ext}(\mathsf{K})$, Proposition 9.6 yields that $\mathsf{K}$ is the class of subreducts of $\mathsf{K}[{L}_{\mathcal{G}}]$. Together with $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$ (which holds because $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$), this entails that there exists $\bm{C}[{L}_{\mathcal{G}}]\in\mathsf{K}[{L}_{\mathcal{G}}]$ such that $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{C}$.
@@
The first equality above holds because $\bm{A}\in\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}]$ and the second is a consequence of the definition of $f_{\bullet}$. The third equality follows from Proposition 8.1 because $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{C}$, and the last from the interpretation of $g_{f}$ in $\mathsf{K}[{L}_{\mathcal{G}}]$. Therefore, $\bm{A}\leqslant\bm{C}[{L}_{\mathcal{G}}]\in\mathsf{K}[{L}_{\mathcal{G}}]$. We conclude that $\mathsf{M}[{L}_{\mathcal{G}_{\bullet}}]\subseteq\mathbb{S}(\mathsf{K}[{L}_{\mathcal{G}}])$, as desired.